A stone is thrown into the air at a 15 m/s at an angle of 30 degree.

a. What is the height reached of the stone?


b. What is the time taken by the stone in the whole travel?


C. What is the ranged reached by the stone?


d. A ball is projected with initial velocity of 18 m/s at an angle of 25 degree. What is the time of flight of the ball?


e.What is the horizontal displacement of the object thrown into the air at a velocity of 30 m/s with an angle of 28 degree?


1
Expert's answer
2021-11-25T10:14:44-0500

a)


"h=\frac{15^2\sin^2{30}}{2(9.8)}=2.9\ m"

b)


"t=2\frac{15\sin{30}}{(9.8)}=1.5\ s"

c)


"R=\frac{15^2\sin{60}}{(9.8)}=20\ m"

d)


"t=2\frac{18\sin{25}}{(9.8)}=1.6\ s"

e)


"R=\frac{30^2\sin{56}}{(9.8)}=76\ m"